OpenStudy (anonymous):

-sqrt2/2 divided by 5pi/2-pi?

OpenStudy (anonymous):

What are we doing here? I'm sorrry.

OpenStudy (anonymous):

\[-\frac{ \sqrt{2} }{ 2 }\div \frac{ 5\pi }{ 2 }-\pi\] if it's clearer that way.

OpenStudy (anonymous):

Ok.. Are we finding the GCF (Greatest Common Factor) or what?

OpenStudy (anonymous):

Well it's a trigonometry question. I'm finding the average rate of change. I just am having trouble with dividing the radical expression by an expression with pi in it.

OpenStudy (anonymous):

Ok no problem I completely understand.

OpenStudy (anonymous):

Do you know what your first step is going to be in this problem?

OpenStudy (anonymous):

Would I multiply top and bottom by pi?

OpenStudy (anonymous):

Or 2?

OpenStudy (anonymous):

pi is correct..Good job!

OpenStudy (anonymous):

Tell me what you get when you do that.

OpenStudy (anonymous):

Take all the time you need. :)

OpenStudy (anonymous):

Okay, thanks. :)

OpenStudy (anonymous):

Absoulutely!

OpenStudy (anonymous):

Hm. I don't think this is right but I got the result as \[\frac{ \sqrt{2} }{ 4\pi }\]

OpenStudy (anonymous):

ok that's a great start actually!!

OpenStudy (anonymous):

Let me show you what I got.

OpenStudy (anonymous):